One leg of a right triangle is 7 cm less than the length of the other leg. The length of the hypotenuse is 13 cm. Find the lengths of the legs.
Answer by mananth(16946)

let one leg be x
the other leg will be x-7
..
right angled triangle
one side ^2 + second side ^2 = hypotenuse ^2
x^2+(x-7)^2=hyp^2
x^2+x^2-14x+49=169
2x^2-14x+120=0
x^2-7x+60=0
x^2-12x+5x+60=0
x(x-12)+5(x-12)=0
(x-12)(x+5)=0
x=12 OR -5.
Ignore the negative value
one leg = 12 cm
the other leg = 12-7=5 cm